OpenStudy (anonymous):

What is the slope of the line which passes through (3, –2) and (–4, –3)?

OpenStudy (lalaly):

\[\underline{slope}=\frac{y_2-y_1}{x_2-x_1} = \frac{-3-(-2)}{-4-3}\]

OpenStudy (anonymous):

rise / run

OpenStudy (asnaseer):